The Taguchi loss function is estimated to be L(x) = 70,000(x – 0.300)2. What are the economic design specifications if the cost of inspection and adjustment is $8.50? Round your answers to three decimal places.

Break-Even Tolerance:_____________

Lower Specification Limit: ______________

Upper Specification Limit:_____________
Answer #1

L(x) = 70000*(x-0.300)^2

8.50 = 70000*(x-0.300)^2

(x-0.300)^2 = 8.5/70000

x-0.300 = +/- 0.011019 (Tolerance Limit)

X = (0.300+0.011019),(0.300- 0.011019)

x = 0.311019, 0.288981

Tolerance = 0.011

LSL = 0.288

USL = 0.311

Note: Please like the answer, if it is helpful to you. Please mention the reason in the comment section, before disliking. Thank You.......!